# Byzantine-resilient Federated Learning With Adaptivity to Data Heterogeneity

### v3

@@ V Experiments
TABLE II: Test accuracy of RAGA, FedAvg, Byrd-SAGA, and RFA with five intensity levels of Byzantine attack on LeNet and MLP learning models. The iteration number $T$ of CIFAR10 and MNIST are 10000 and 1000 respectively.
Dataset CIFAR10 MNIST
Model LeNet MLP LeNet MLP
Level of Byzantine Attacks ($1-C_{\alpha}$) 0 0.1 0.2 0.3 0.4 0 0.1 0.2 0.3 0.4 0 0.1 0.2 0.3 0.4 0 0.1 0.2 0.3 0.4
RAGA (%) 79.25 79.09 78.18 77.13 77.14 58.46 59.04 57.74 58.17 56.59 96.53 96.57 96.50 96.39 96.27 92.90 92.91 92.77 92.84 92.69
FedAvg (%) 80.02 10.00 10.00 10.00 10.00 59.76 10.00 10.00 10.00 10.00 96.92 11.35 11.35 11.35 10.32 93.07 11.35 10.28 10.28 9.58
Byrd-SAGA (%) 72.13 72.35 70.46 72.33 71.98 56.10 55.94 55.57 55.85 55.50 93.01 93.06 92.91 92.57 92.90 90.99 90.91 90.93 90.69 90.78
RFA (%) 78.07 77.42 77.14 76.60 75.52 58.86 58.18 57.79 57.59 56.44 96.01 95.96 95.91 95.87 95.86 92.49 92.48 92.52 92.35 92.34
Figure 2: Text accuracy and training loss of LeNet with CIFAR10 dataset for RAGA, FedAvg, Byrd-SAGA, and RFA. The intensity levels of Byzantine attack ($1-C_{\alpha}$) are 0, 0.2, 0.4 from left to right.
Figure 3: Text accuracy and training loss of MLP with CIFAR10 dataset for RAGA, FedAvg, Byrd-SAGA, and RFA. The intensity levels of Byzantine attack ($1-C_{\alpha}$) are 0, 0.2, 0.4 from left to right.
Figure 4: Text accuracy and training loss of LeNet with MNIST dataset for RAGA, FedAvg, Byrd-SAGA, and RFA. The intensity levels of Byzantine attack ($1-C_{\alpha}$) are 0, 0.2, 0.4 from left to right.
Figure 5: Text accuracy and training loss of MLP with MNIST dataset for RAGA, FedAvg, Byrd-SAGA, and RFA. The intensity levels of Byzantine attack ($1-C_{\alpha}$) are 0, 0.2, 0.4 from left to right.
@@ V-A Setup
To carry out experiments, we set up a machine learning environment in PyTorch 2.1.0 on Ubuntu 20.04, powered by two 3090 GPUs.
V-A1 Datasets and Models
We conduct experiments by leveraging two datasets: MNIST and CIFAR10. These two datasets are widely used for learning tasks and are introduced as follows:
@@ V-A1 Datasets and Models
  • MNIST: The MNIST dataset includes a training set and a test set. The training set contains 60000 samples and the test set contains 10000 samples, each sample of which is a 28 $\times$ 28 pixel grayscale image.
@@
  • CIFAR10: The CIFAR10 dataset is composed of a training set and a test set. The training set contains 50000 samples and the test set contains 10000 samples, where each sample is a 32 $\times$ 32 color image.
@@
For these experiments, we split the above two training sets into $M$ non-IID training sets, which is realized by letting the label of data samples to conform to Dirichlet distribution. The extent of non-IID can be adjusted by tuning the concentration parameter of Dirichlet distribution.
@@
For training problems with non-convex and strongly-convex loss functions, we adopt LeNet and Multilayer Perceptron (MLP) model, respectively. The introduction of these two models are as follows:
@@
  • LeNet: The LeNet model is one of the earliest published convolutional neural networks. For the experiments, we are going to train a LeNet model with two convolutional layers, two pooling layers, and one fully connected layer like in [35]. Cross-entropy function is taken as the training loss.
@@
  • MLP: The MLP model is a machine learning model based on Feedforward Neural Network that can achieve high-level feature extraction and classification. To promise strong convexity for the loss function, we configure the MLP model to be with only fully connected layers like in [36] and take the cross-entropy function as the training loss.
V-A2 Hyperparameters
We assume $M=50$ and the batchsize of all the experiments is 512. The error tolerance $\epsilon$ for numerically working out geometric median is $\epsilon=1\times 10^{-5}$. The concentration parameter of Dirichlet distribution is 0.6. In default, the iteration number $T=1000$ for the MNIST dataset, and $T=10000$ for the CIFAR10 dataset. The round number of local updates $K^{t}$ is 3 for RFA and RAGA. The Byzantine attacks are selected as the Gaussian attack, which obeys standard normal distribution. $1-C_{\alpha}$ are used to indicate the intensity level of Byzantine attacks. In terms of the learning rate, we adopt default setup as shown in corresponding papers for baseline methods. For RAGA, the learning rate $\eta^{t,k}_{m}=\eta^{t}={T}/{(100t+10T)}$.
V-A3 Baselines
@@
The convergence performance of four algorithms: RAGA, FedAvg [17], Byrd-SAGA [28], and RFA [27] are compared. For the latter three algorithms, FedAvg is renowned in traditional FL but does not concern Byzantine attack and can be taken as a performance metric when there is no Byzantine attack, Byrd-SAGA only executes one local updating in each round of iteration, and RFA selects the trimmed mean of the model parameters over multiple local updating as the uploaded vector. It is worthy to note that there is another Byzantine-resilient FL algorithm in related works: BROADCAST [29]. We do not show the performance of it because BROADCAST is very similar with Byrd-SAGA and only differs from Byrd-SAGA by uploading vector from each user to the central server in a compressive, rather than lossless way like in Byrd-SAGA.
V-A4 Performance Metrics
To assess the convergence performance, we measure the test accuracy, the convergence speed and the training loss. An algorithm would be better if its test accuracy is higher, convergence speed is faster, or training loss is lower.
@@ V-B Convergence Performance
We start from Table II, which shows the test accuracy of these four algorithms under different intensity levels of Byzantine attack for various datasets and learning models. From Table II, we see that RAGA can mostly achieve the highest test accuracy under Byzantine attacks. Even there is no Byzantine attack, in which case FedAvg performs best, RAGA ranks the second in most cases and is very close to FedAvg. Additionally, as the intensity level of Byzantine attack grows from 0 to 0.4, in which case nearly half of training data of the whole system is corrupted by Byzantine users, we see RAGA degrades little, which verifies the robustness of RAGA to Byzantine attacks.
Fig. 2 and Fig. 3 illustrate the convergence performance of the four comparing algorithms versus iteration number, by means of test accuracy and training loss, in the training task of LeNet and MLP respectively, for CIFAR10 dataset. From Fig. 2 and Fig. 3, we can observe that both RAGA and FedAvg act similarly and perform better than other algorithms when there is no Byzantine attack. With the existence of Byzantine attack, FedAvg performs badly as it does not take Byzantine attack into account and RAGA becomes the one that outperforms all the other algorithms, both in terms of test accuracy and training loss.
Fig. 4 and Fig. 5 repeat the experiment as shown in Fig. 2 and Fig. 3 but for a different dataset, i.e., MNIST dataset. From Fig. 4 and Fig. 5, when there is no Byzantine attack, we can see that both RAGA and FedAvg have the best convergence performance but RAGA is more stable than FedAvg for the training task of LeNet model. As Byzantine attack comes up, FedAvg is very sensitive to the Byzantine attack and degrades a lot. In contrast, RAGA is always the best one compared with all the other algorithms, no matter how strong the Byzantine attack is.
In summary, through the above numerical experiments, it can be seen that the RAGA not only defeats its opponents at different levels of Byzantine attack, but also acts well and has a comparable performance with FedAvg that performs best when there is no Byzantine attack. It is also worthy to mention that RAGA achieves a nearly unchanged convergence performance no matter how strong the Byzantine attack is. Last but not least, all the above results are obtained based on non-IID datasets.

### v4

@@ V Experiments
TABLE II: The maximum test accuracy(%) of 500 iteration number of RAGA, FedAvg, Median, Byrd-SAGA, RFA and MCA with different types, ratios of Byzantine attack, and concentration parameter on MNIST dataset and LeNet model.
Model Attack Name No Attack Gaussian Attack Sign-flip Attack LIE Attack
$\bar{C}_{\alpha}$ 0 0.1 0.2 0.3 0.4 0.1 0.2 0.3 0.4 0.1 0.2 0.3 0.4
$\phi$ Algorithms
LeNet 0.6 RAGA 97.89 97.78 97.79 97.82 97.75 97.72 97.76 97.69 97.25 97.79 97.82 97.80 97.71
FedAvg 95.54 10.28 11.35 11.35 11.35 94.03 9.80 9.82 9.80 95.10 94.96 93.66 92.63
Median 94.41 94.31 94.22 94.53 94.18 94.12 94.34 93.69 93.45 94.33 94.35 94.60 94.22
Byrd-SAGA 91.70 91.66 91.61 91.77 91.26 91.23 91.54 91.13 89.95 91.64 91.66 91.87 91.30
RFA 95.70 95.52 95.53 95.74 95.41 95.39 95.05 95.23 94.85 95.54 95.48 95.51 95.26
MCA 91.82 91.89 91.88 92.16 91.77 90.77 19.02 83.10 18.75 91.86 91.94 92.14 91.76
0.2 RAGA 97.62 97.47 97.67 97.50 97.30 97.42 97.16 97.28 97.38 97.75 97.64 97.60 97.28
FedAvg 95.13 9.80 11.35 10.32 9.80 10.10 10.10 10.09 9.80 94.81 94.38 93.74 92.09
Median 90.81 91.88 91.43 91.03 91.91 89.85 89.05 90.28 88.28 91.68 91.20 90.90 92.00
Byrd-SAGA 90.36 90.76 90.24 89.81 89.77 89.51 87.95 89.28 88.29 90.90 90.35 89.69 89.82
RFA 95.09 95.23 95.06 94.99 94.61 93.67 93.67 93.87 94.32 95.15 95.07 94.74 94.28
MCA 90.96 91.44 91.04 90.85 90.72 90.04 85.89 36.78 9.80 91.50 90.99 90.67 90.79
(a) LeNet and $\bar{C}_{\alpha}=0.2$.
(b) LeNet and $\bar{C}_{\alpha}=0.4$.
(c) MLP and $\bar{C}_{\alpha}=0.2$.
(d) MLP and $\bar{C}_{\alpha}=0.4$.
Figure 2: Test accuracy of RAGA and baselines on $\phi=0.6$, Sign-flip attack, and MNIST dataset.
@@ V-A Setup
To carry out experiments, we set up a machine learning environment in PyTorch 2.3.1 on Ubuntu 20.04, powered by two 3090 GPUs and two Intel Xeon Gold 6226R CPUs.
Datasets:
@@
  • MNIST: MNIST dataset includes a training set and a test set. The training set contains 60000 samples and the test set contains 10000 samples, each sample of which is a 28 × 28 pixel grayscale image.
@@
  • CIFAR10: The CIFAR10 dataset includes a training set and a test set. The training set contains 50,000 samples, and the test set contains 10,000 samples, each of which is a 32 × 32 pixel color image.
  • CIFAR100: The CIFAR100 dataset includes a training set and a test set. The training set contains 50,000 samples, and the test set contains 10,000 samples, each of which is a 32 × 32 pixel color image.
TABLE III: The maximum test accuracy(%) of 500 iteration number of RAGA, FedAvg, Median, Byrd-SAGA, RFA and MCA with different types, ratios of Byzantine attack, and concentration parameter on MNIST dataset and MLP model.
Model Attack Name No Attack Gaussian Attack Sign-flip Attack LIE Attack
$\bar{C}_{\alpha}$ 0 0.1 0.2 0.3 0.4 0.1 0.2 0.3 0.4 0.1 0.2 0.3 0.4
$\phi$ Algorithms
MLP 0.6 RAGA 95.23 95.10 95.06 95.01 94.98 94.84 95.00 94.23 95.01 95.08 95.16 95.02 94.92
FedAvg 92.09 23.10 24.69 25.17 9.84 89.41 10.10 9.80 10.32 91.92 90.44 87.98 85.85
Median 91.48 91.38 91.63 91.29 91.24 91.23 91.32 90.94 91.55 91.39 91.46 91.26 91.27
Byrd-SAGA 90.64 90.69 90.70 90.49 90.63 90.46 90.62 89.93 90.62 90.64 90.69 90.59 90.57
RFA 92.38 92.31 92.41 92.04 92.27 92.10 91.86 90.81 92.33 92.36 92.46 92.05 92.18
MCA 90.62 90.66 90.77 90.52 90.63 90.36 87.86 60.77 10.32 90.66 90.69 90.56 90.70
0.2 RAGA 94.53 94.45 94.69 94.16 94.32 94.09 94.43 93.84 94.72 94.35 94.70 94.16 94.38
FedAvg 91.84 21.48 19.66 30.05 12.34 10.10 11.98 9.80 10.09 91.48 90.27 87.50 85.41
Median 89.54 89.16 89.74 89.04 89.37 88.65 89.70 87.14 89.83 89.22 89.70 89.18 89.51
Byrd-SAGA 90.24 90.12 90.08 89.96 89.76 89.86 90.08 88.40 90.06 90.04 90.08 89.92 89.85
RFA 92.01 91.88 91.92 91.65 91.66 91.72 91.91 91.23 92.04 91.89 91.95 91.62 91.49
MCA 90.32 90.21 90.40 90.23 89.99 89.63 81.01 94.81 10.28 90.29 90.32 90.21 89.91
@@
We split the above three datasets into $M$ non-IID training sets, which is realized by letting the label of data samples to conform to Dirichlet distribution. The extent of non-IID can be adjusted by tuning the concentration parameter $\phi$ of Dirichlet distribution.
@@
Models: We adopt LeNet, Multilayer Perceptron (MLP), and VGG16 model. The introduction of these three models are as follows:
@@
  • LeNet: The LeNet model is one of the earliest published convolutional neural networks. For the experiments, like [39], we are going to train a LeNet model with two convolutional layers (both kernel size are 5 and the out channel of first one is 6 and second one is 16), two pooling layers (both kernel size are 2 and stride are 2), and three fully connected layer (the first one is (16 × 4 × 4, 120), the second is (120, 60) and the last is (60, 10)) for MNIST dataset. And for CIFAR10 dataset, we are going to train a LeNet model with two convolutional layers (both kernel size are 5 and the out channel are 64), two pooling layers (both kernel size are 2 and stride are 2), and three fully connected layer (the first one is (64 × 5 × 5, 384), the second is (384, 192) and the last is (192, 10)). Cross-entropy function is taken as the training loss.
@@
  • MLP: The MLP model is a machine learning model based on Feedforward Neural Network that can achieve high-level feature extraction and classification. We configure the MLP model to be with three connected layers (the first one is (28 × 28, 200), the second is (200, 100) and the last is (100, 10)) like [40]. And for CIFAR10 dataset, we configure the MLP model to be with three connected layers (the first one is (3× 32 × 32, 200), the second is (200, 200) and the last is (200, 10)). And also cross-entropy function is taken as the training loss.
  • VGG16: The VGG16 model, originally developed by Simonyan and Zisserman from the Visual Geometry Group, represents a seminal 16-layer convolutional neural network architecture comprising 13 convolutional layers and 3 fully-connected (FC) layers. Each convolutional stage utilizes cascaded 3×3 kernels with stride 1 and ReLU activation, interspersed with 2×2 max-pooling operations that halve spatial resolution while preserving depth. The fully-connected hierarchy consists of two 4,096-unit hidden layers (FC1-2) followed by a 1,000-class output layer (FC3), totaling 138M trainable parameters. For CIFAR100 adaptation, we implemented fine-tuning to adapt to this dataset.
(a) LeNet and $\bar{C}_{\alpha}=0.2$.
(b) LeNet and $\bar{C}_{\alpha}=0.4$.
(c) MLP and $\bar{C}_{\alpha}=0.2$.
(d) MLP and $\bar{C}_{\alpha}=0.4$.
Figure 3: Test accuracy of RAGA and baselines on $\phi=0.6$, Sign-flip attack, and CIFAR10 dataset.
Byzantine attacks: The intensity level of Byzantine attack $\bar{C}_{\alpha}=1-C_{\alpha}$ is set to 0, 0.1, 0.2, 0.3, 0.4 and we select three types of Byzantine attack, which are introduced as follows:
  • Gaussian attack: All Byzantine attacks are selected as the Gaussian attack, which obeys $\mathcal{N}(0,90)$.
  • Sign-flip attack: All Byzantine clients upload $-3*\sum_{m\in\mathcal{M}\setminus\mathcal{B}}\bm{z}_{m}^{t}$ to the central server on iteration number $t$.
  • LIE attack [32]: LIE attack adds small amounts of noise to each dimension of the benign gradients. The noise is controlled by a coefficient $c$, which enables the attack to evade detection by robust aggregation methods while negatively impacting the global model. Specifically, the attacker calculates the mean $\rho$ and standard deviation $\nu$ of the parameters submitted by honest users, calculates the coefficient $c$ based on the total number of honest and malicious clients, and finally computes the malicious update as $\rho$ + $c\nu$. We set $c$ to 0.7.
Hyperparameters: We assume $M=50$ and the batchsize of all the experiments is 32. The error tolerance $\epsilon$ for numerically working out geometric median and MCA are $\epsilon=1\times 10^{-5}$. The concentration parameter $\phi$ is set to 0.6, 0,4 and 0.2. The iteration number $T=500$ for both MNIST and CIFAR10 dataset, while $T=100$ for CIFAR100 dataset. In default, the round number of local updates $K^{t}=3$ for all algorithms on LeNet and MLP model, while $K^{t}=1$ on VGG16 model. In terms of the learning rate, we adopt default setup as shown in corresponding papers for baseline methods. For RAGA, the learning rate $\eta^{t,k}_{m}=\eta^{t}=\frac{K^{t}}{\sqrt{5}\sqrt{t+5}}=\frac{K^{t}}{5\sqrt{0.2t+1}}$ for LeNet and MLP models, while $\eta^{t,k}_{m}=\eta^{t}=\frac{4*10^{-2}K^{t}}{\sqrt{t+100}}=\frac{4*10^{-3}K^{t}}{\sqrt{0.01t+1}}$ for VGG16 model. This learning rate configuration is motivated by Remark 1, which is derived from Theorem 1 and tailored for non-convex loss functions, covering the experimental models including LeNet, MLP, and VGG16. It is worth noting that the learning rate used in practice slightly deviates from that suggested in Remark 1 due to implementation considerations. Specifically, we set $\delta=1/6$ and $c=5$ for LeNet and MLP, and $\delta=1/6$ and $c=100$ for VGG16. Since the exact smoothness constant $L$ is unknown, we adopt heuristic coefficients, e.g., $\frac{K^{t}}{\sqrt{5}}$ for LeNet/MLP and $4*10^{-2}K^{t}$ for VGG16, to ensure the learning rate remains sufficiently small for convergence.
@@
Baselines: The convergence performance of four algorithms: RAGA, FedAvg [16], Median [15], Byrd-SAGA [28], RFA [27], MCA [26] are compared. For the latter five algorithms, FedAvg is renowned in traditional FL but does not concern Byzantine attack and can be taken as a performance metric when there is no Byzantine attack, Median takes the coordinate-wise median of the uploaded vectors to update the global model parameter, Byrd-SAGA only executes one local updating in each round of iteration, RFA selects the trimmed mean of the model parameters over multiple local updating as the uploaded vector, and MCA utilizes its the maximum correntropy of the model parameters’ distribution characteristics and compute the central value to update the global model parameter. It is worthy to note that there is another Byzantine-resilient FL algorithm in related works: BROADCAST [29]. We do not show the performance of it because BROADCAST is very similar with Byrd-SAGA and only differs from Byrd-SAGA by uploading vector from each user to the central server in a compressive, rather than lossless way like in Byrd-SAGA.
@@ V-B Convergence Performance
(a) LeNet and Gaussian attack.
(b) LeNet and LIE attack.
(c) MLP and Gaussian attack.
(d) MLP and LIE attack.
Figure 4: The maximum test accuracy(%) of 500 iteration number of RAGA and baselines on $\phi=0.4$ and MNIST dataset.
TABLE IV: The maximum test accuracy(%) of 500 iteration number of RAGA, FedAvg, Median, Byrd-SAGA, RFA and MCA with different types, and ratios of Byzantine attack on CIFAR10 dataset and two learning models.
Model Attack Name No Attack Gaussian Attack Sign-flip Attack LIE Attack
$\bar{C}_{\alpha}$ 0 0.1 0.2 0.3 0.4 0.1 0.2 0.3 0.4 0.1 0.2 0.3 0.4
$\phi$ Algorithms
LeNet 0.4 RAGA 42.92 43.28 41.68 42.81 41.89 41.58 41.40 41.03 42.29 42.91 40.20 42.12 41.42
FedAvg 39.45 10.91 10.00 10.00 10.00 10.00 10.00 10.00 10.00 41.37 35.38 25.40 15.89
Median 32.17 31.70 30.83 31.59 30.29 30.09 30.54 28.00 30.19 31.50 30.83 32.25 31.25
Byrd-SAGA 34.81 34.60 33.86 33.92 34.27 33.49 33.66 29.37 34.21 35.20 34.19 33.89 34.64
RFA 41.85 41.64 41.65 41.42 41.27 40.92 41.61 37.20 41.65 41.78 41.32 41.21 40.88
MCA 35.10 34.83 34.26 34.23 34.69 32.92 18.01 10.00 10.00 34.42 34.21 34.64 34.53
MLP 0.4 RAGA 44.21 44.32 44.06 44.29 43.13 43.63 43.71 44.02 42.88 44.11 43.93 44.65 43.20
FedAvg 40.13 10.86 10.00 12.03 13.84 12.32 14.05 14.46 10.00 34.38 30.80 25.30 23.83
Median 37.66 37.34 37.48 37.62 36.51 36.77 37.06 37.12 36.23 37.56 37.42 37.76 36.54
Byrd-SAGA 37.38 36.72 36.93 36.85 36.01 36.34 37.03 36.52 35.99 36.78 37.03 36.89 36.18
RFA 40.53 40.41 40.26 40.94 39.85 37.91 40.28 39.61 40.08 40.69 40.51 40.72 39.97
MCA 37.18 36.87 36.94 36.99 36.36 36.47 14.68 15.40 10.00 36.63 37.18 37.00 35.97
We begin with Table II, III, IV and V, which presents the maximum test accuracy of RAGA and baselines under varying levels of Byzantine attack intensity $\bar{C}_{\alpha}$, across different data heterogeneity concentration parameters $\phi$, and for two learning models. Fig 2 and 3 illustrates the convergence of RAGA and the baselines under Sign-flip attack and data heterogeneity concentration parameters $\phi=0.6$. Fig 4, 5 and 6 present the maximum test accuracy of RAGA and the baseline methods under various Byzantine ratios, excluding the results reported in Table II, III and IV. Fig 7, 8 present the test accuracy of RAGA with different data heterogeneity parameters under various Byzantine ratios. And Fig 9 and 10 illustrates convergence of RAGA under different local epoch $K$ on Gaussian attack with $\bar{C}_{\alpha}=0.2$. Following this, we will discuss the performance comparison between RAGA and the baselines in different scenarios.
(a) LeNet and Gaussian attack.
(b) LeNet and LIE attack.
(c) MLP and Gaussian attack.
(d) MLP and LIE attack.
Figure 5: The maximum test accuracy(%) of 500 iteration number of RAGA and baselines on $\phi=0.6$ and CIFAR10 dataset.
No Byzantine attack: For the easy dataset (MNIST), Table II, III and Fig. 4 show that RAGA outperforms all other methods across a range of data heterogeneity concentration parameters, $\phi$, and two different learning models. Specifically, compared to the baselines, RAGA increases test accuracy by 2.19% and 2.85% for the LeNet and MLP models, respectively, with $\phi=0.6$, and by 2.49% and 2.52% for the LeNet and MLP models with $\phi=0.2$. For the CIFAR10 dataset, Table IV and Fig. 5 demonstrate that RAGA also achieves the highest performance across different data heterogeneity concentrations, particularly for $\phi=0.4$ with the LeNet and MLP models. In comparison to the baselines, RAGA increases test accuracy by 1.07% and 3.68% for the LeNet and MLP models, respectively, at $\phi=0.4$, and consistently shows the highest accuracy at $\phi=0.6$. As presented in Table V, when applied to the CIFAR100 dataset, RAGA demonstrates a notable improvement in test accuracy, achieving a 3.89% increase compared to the robust baselines. In contrast, the performance gap between RAGA and FedAvg is minimal, narrowing to just 0.1%.
TABLE V: The maximum test accuracy(%) of 100 iteration number of RAGA, FedAvg, Median, Byrd-SAGA, RFA and MCA with different types, and ratios of Byzantine attack on CIFAR100 dataset and VGG16 model.
Model Attack Name No Attack Gaussian Attack Sign-flip Attack LIE Attack
$\bar{C}_{\alpha}$ 0 0.1 0.2 0.3 0.4 0.1 0.2 0.3 0.4 0.1 0.2 0.3 0.4
$\phi$ Algorithms
VGG16 0.6 RAGA 54.57 54.85 54.42 53.73 54.03 54.19 54.48 53.66 54.40 54.61 54.55 54.62 53.71
FedAvg 54.67 29.59 20.73 14.06 10.02 17.45 1.00 1.00 21.88 53.38 52.03 50.71 49.86
Median 49.51 49.89 49.50 48.79 49.81 49.17 49.15 47.42 48.01 49.39 48.97 48.80 48.75
SAGA 50.50 50.78 50.19 49.11 48.64 49.88 50.30 49.60 49.38 50.47 50.05 49.66 48.91
RFA 50.68 50.80 50.22 49.01 48.57 50.62 50.29 50.09 49.55 50.42 50.04 49.73 48.92
MCA 50.48 50.64 50.17 49.21 48.66 49.41 7.32 1.00 1.00 50.43 50.00 49.63 48.88
(a) Gaussian attack.
(b) Sign-flip attack.
(c) LIE attack.
Figure 6: The maximum test accuracy (%) of 100 iteration number of RAGA and baselines on $\phi=0.6$ and CIFAR100 dataset.
Gaussian attack: First, we evaluate the performance of RAGA and the baseline methods on the MNIST dataset. As shown in Table II and III, RAGA achieves the highest test accuracy among all FL robust algorithms across different Byzantine attack ratios, two learning models, and two distinct data heterogeneity concentration parameters. For the LeNet model, RAGA shows an improvement in test accuracy by 2.08% to 2.34% for the data heterogeneity concentration parameter $\phi=0.6$, and by 2.24% to 2.69% for $\phi=0.2$, across four different Byzantine ratios ($\bar{C}_{\alpha}=0.1,0.2,0.3,0.4$). For the MLP model, RAGA improves test accuracy by 2.65% to 2.97% for $\phi=0.6$, and by 2.57% to 2.77% for $\phi=0.2$, across the same four Byzantine ratios. Additionally, as shown in Fig. 4, RAGA outperforms five other FL robust algorithms across various Byzantine attack ratios, regardless of the learning model type, when the data heterogeneity concentration parameter is set to $\phi=0.4$. Next, we analyze the performance of RAGA and the baseline methods on the more complex CIFAR-10 dataset. With the data heterogeneity concentration parameter $\phi=0.4$, RAGA consistently outperforms all baseline methods in terms of test accuracy, as shown in Table IV. Specifically, for the LeNet model, RAGA achieves improvements by 0.03% to 1.64%, while for the MLP model, improvements by 3.25% to 3.91%, across four different Byzantine ratios ($\bar{C}_{\alpha}=0.1,0.2,0.3,0.4$). Furthermore, RAGA consistently performs better than all baseline methods based on Fig. 5, irrespective of whether the LeNet or MLP model is used, when the data heterogeneity concentration parameter is set to $\phi=0.6$. For the CIFAR100 dataset, Table V demonstrates that RAGA achieves a maximum accuracy improvement of 4.05% to 4.52% across four distinct Byzantine ratios when compared to baseline methods. Additionally, as shown in Fig. 6, RAGA consistently outperforms all baselines on the VGG16 model under Gaussian attacks.
(a) MNIST and Gaussian attack.
(b) MNIST and Sign-flip attack.
(c) MNIST and LIE attack.
(d) CIFAR10 and Gaussian attack.
(e) CIFAR10 and Sign-flip attack.
(f) CIFAR10 and LIE attack.
Figure 7: Test accuracy of RAGA with different data heterogeneity parameters on $\bar{C}_{\alpha}=0.4$.
(a) Sign-flip attack with $\bar{C}_{\alpha}=0.1$.
(b) LIE attack with $\bar{C}_{\alpha}=0.1$.
(c) Gaussian attack with $\bar{C}_{\alpha}=0.4$.
(d) LIE attack with $\bar{C}_{\alpha}=0.4$.
Figure 8: Test accuracy of RAGA with different data heterogeneity parameters on CIFAR100 dataset.
Sign-flip attack: First, for the relatively simple dataset (MNIST), as shown in Table II and III, all FL robust methods, with the exception of MCA, effectively mitigate the sign-flip attack across various Byzantine ratios. In contrast, MCA fails to defend against this attack at higher Byzantine ratios. For the LeNet model, Table II and III show that RAGA outperforms all other methods in terms of test accuracy, improving by 2.33% to 2.71% with the data heterogeneity concentration parameter of $\phi=0.6$, and by 3.06% to 3.75% with $\phi=0.2$, across four different Byzantine ratios. For the MLP model, RAGA also achieves the highest test accuracy among all baseline methods, improving by 2.68% to 3.42% for $\phi=0.6$, and by -0.97% to 2.68% for $\phi=0.2$, across the same four Byzantine ratios. Then, for the more complex dataset (CIFAR10), as shown in Table IV, similar to the results on the MNIST dataset, all FL robust methods, except MCA, effectively mitigate the sign-flip attack across various Byzantine ratios. With the data heterogeneity concentration parameter of $\phi=0.4$, RAGA performs the best in most cases (except for the LeNet model with $\bar{C}_{\alpha}=0.2$), with test accuracy improvements of -0.21% to 4.01% for the LeNet model, and 2.8% to 5.72% for the MLP model, across the four Byzantine ratios. Additionally, as shown in Fig. 2 and 3, RAGA achieves the fastest convergence speed among all baseline methods with the same local epoch $K^{t}=3$ and data heterogeneity concentration parameter $\phi=0.6$, regardless of the learning model type or the Byzantine ratio. According to Table V, for the CIFAR100 dataset, RAGA consistently boosts test accuracy by 3.57% to 4.85% across four distinct Byzantine ratios compared to baselines when $\phi=0.6$. Fig. 6 further illustrates that RAGA outperforms all baselines across the evaluated conditions. Regarding the FedAvg result with $C_{\alpha}=0.4$ that exhibits a seemingly ”maximum” accuracy, this outcome should not be attributed to statistical abnormality. Instead, the extremely high loss value of 856667.4125 indicates that the model likely converged to a distant saddle point in the loss landscape.
LIE attack: Using the simple MNIST dataset, RAGA achieves the best performance among six baseline methods across four Byzantine ratios. As shown in Table II and III, for the LeNet model, RAGA improves test accuracy by 2.25% to 2.45% when the data heterogeneity concentration parameter is $\phi=0.6$, and by 2.57% to 3.00% when $\phi=0.2$, across four different Byzantine ratios. Similarly, for the MLP model, RAGA outperforms all baseline methods, improving test accuracy by 2.7% to 2.97% when the data heterogeneity concentration parameter is $\phi=0.6$, and by 2.46% to 2.89% when $\phi=0.2$. Furthermore, as shown in Fig. 4, with a data heterogeneity concentration parameter of $\phi=0.4$, RAGA achieves the best test accuracy performance among five FL Byzantine-robust algorithms, regardless of the type of learning model. Additionally, using the more complex CIFAR-10 dataset, RAGA outperforms most baseline methods (with the exception of the LeNet model at a Byzantine ratio of $\bar{C}_{\alpha}=0.2$). With the data heterogeneity concentration parameter of $\phi=0.4$, RAGA enhances test accuracy by -1.12% to 1.13% on the LeNet model, and by 3.23% to 3.93% on the MLP model, across four Byzantine ratios, compared to other FL Byzantine-robust algorithms. Finally, as shown in Fig. 5, under the LIE attack, RAGA also outperforms all six baseline methods with the data heterogeneity concentration parameter $\phi=0.6$. Regarding the CIFAR100 dataset, RAGA demonstrates superior performance against LIE attacks, achieving improvements in test accuracy ranging from 1.23% to 3.91% across four distinct Byzantine ratios when compared to baseline methods, as shown in Table V. Additionally, Fig. 6 confirms that RAGA establishes state-of-the-art (SoTA) performance under these experimental settings, outperforming all baseline approaches evaluated.
Data Heterogeity: We first analyze the convergence performance of RAGA on the MNIST dataset. As shown in Table II and III, varying data heterogeneity parameters has minimal impact on the test accuracy of RAGA, with the maximum difference being only 0.66%, as further confirmed by Fig. 7. In contrast, for the more complex CIFAR10 dataset, the effect of different data heterogeneity parameters on RAGA’s test accuracy is more pronounced, as depicted in Fig. 7. Additionally, we observe that the MLP model is more robust to data heterogeneity than the LeNet model, particularly in the presence of the Sign-flip attack. For the CIFAR100 dataset, as depicted in Fig. 8, the experimental results on the VGG16 model demonstrate that the impact of attack type on accuracy surpasses that of data heterogeneity. This distinction is notably evident when comparing the Sign-flip and LIE attacks, where differing attack strategies produce more significant effects on model accuracy than variations in data distribution heterogeneity.
(a) MNIST, and LeNet.
(b) MNIST, and MLP.
(c) CIFAR10, and LeNet.
(d) CIFAR10, and MLP.
Figure 9: Test accuracy of RAGA with different local epoch on Gaussian attack on $\bar{C}_{\alpha}=0.2$ with $\phi=0.6$.
(a) MNIST, and LeNet.
(b) MNIST, and MLP.
(c) CIFAR10, and LeNet.
(d) CIFAR10, and MLP.
Figure 10: Test accuracy of RAGA with different local epoch on Gaussian attack on $\bar{C}_{\alpha}=0.2$ with $\phi=0.2$.
Local update: From Fig. 9 and 10, it is clear that a larger local update size $K$ does not necessarily result in better performance, which is consistent with our theoretical analysis, irrespective of the learning model or dataset. With the LeNet model, the convergence curve exhibits significant fluctuations when $K=1$, leading to suboptimal convergence performance. When $K=5$, although the fluctuations are reduced, the increased number of training rounds adversely affects the algorithm’s performance. In contrast, for $K=2$ or $K=3$, the choice of $K$ demonstrates distinct advantages across both datasets and learning models. These experimental findings are in agreement with our theoretical proof.
In summary, the above numerical experiments demonstrate that RAGA not only outperforms its competitors across different ratios and types of Byzantine attacks but also performs comparably to FedAvg, which excels when there is no Byzantine attack. It is also noteworthy that the RAGA algorithm can converge under various types of Byzantine attacks. Lastly, all these results are based on non-IID datasets with three different data heterogeneity concentration parameters.